# The spectrum of Dirichlet-to-Neumann maps for radial conductivities

### v1

@@ Abstract.
The problem of characterizing sequences of real numbers that arise as spectra of Dirichlet-to-Neumann (DtN) maps for elliptic operators has attracted considerable attention over the past fifty years. In this article, we address this question in the simple setting of DtN maps associated with a rotation-invariant elliptic operator $\operatorname{div}(\gamma\nabla\cdot)$ in the ball in Euclidean space. We show that the spectrum of such a DtN operator can be expressed as a universal term, determined solely by the boundary values of the conductivity $\gamma$, plus a sequence of Hausdorff moments of an integrable function, which we call the Born approximation of $\gamma$. We also show that this object is locally determined from the boundary from the corresponding values of the conductivity, a property that implies a local uniqueness result for the Calderón Problem in this setting. We also give a stability result: the functional mapping the Born approximation to its conductivity is Hölder stable in suitable Sobolev spaces. Finally, in order to refine the characterization of the Born approximation, we analyze its regularity properties and their dependence on the conductivity.
@@ 1. Introduction
where $\partial_{\nu}$ is the outward normal unit vector field on $\partial\Omega$ and $u$ is the solution to the conductivity equation
@@
Since constants are always solutions to (1.2), one always has $\lambda_{0}[\gamma]=0$. We will show that the rest of the eigenvalues of a radial DtN map have a very rigid structure: they can be expressed in terms of moments of radial functions. For $h\in L^{1}(\mathbb{B}^{d})$ define
(1.3) $\sigma_{k}[h]:=\frac{1}{|\mathbb{S}^{d-1}|}\int_{\mathbb{B}^{d}}h(x)|x|^{2k}\,dx,\qquad k\in\mathbb{N}_{0},$
@@
(1.4) $\lambda_{k}[\gamma]=\gamma|_{\mathbb{S}^{d-1}}k-\frac{1}{2}\partial_{\nu}\gamma|_{\mathbb{S}^{d-1}}+\sigma_{k}[\Delta\gamma^{\mathrm{B}}],\qquad\forall k\in\mathbb{N},$
@@
(1.5) $\gamma^{\mathrm{B}}|_{\mathbb{S}_{d-1}}=\gamma|_{\mathbb{S}_{d-1}},\qquad\qquad\partial_{\nu}\gamma^{\mathrm{B}}|_{\mathbb{S}_{d-1}}=\partial_{\nu}\gamma|_{\mathbb{S}_{d-1}}.$
@@
    (1.6) $\sigma_{k}[\gamma^{\mathrm{B}}]=\frac{\lambda_{k}[\gamma]}{2(k+1)(k+1+\nu_{d})},\qquad k\in\mathbb{N}_{0},\qquad\nu_{d}:=\frac{d-2}{2}.$
@@
$\sigma_{k}[\Delta\gamma^{\mathrm{B}}]=\int_{0}^{1}h_{\gamma}(r)r^{k}dr,\qquad k\in\mathbb{N}.$
@@
The function $\gamma^{\mathrm{B}}$ appearing in ˜1 is called the Born approximation of $\gamma$. It is a purely spectral object, in the sense that it is completely determined by $\operatorname{\mathrm{Spec}}(\Lambda_{\gamma})$ as (1.7) shows. However, it is local in the sense that in any annular neighborhood of the boundary $\partial\mathbb{B}^{d}$, it is uniquely determined by the values of the conductivity in that same neighborhood.
@@
These two results refine, in the radial case, structural results for the spectra of DtN maps corresponding to general, smooth, non-radial conductivities obtained, for instance, in [LSA21, Roz78] for the two-dimensional case, or in [PEUVB24] for the ball in Euclidean space, which are based on a detailed analysis of (pseudodifferential or Berezin) symbols of DtN maps. In particular, ˜1 gives a non-trivial description of the spectrum even in the case in which the conductivity $\gamma$ is constant in a neighborhood of the boundary $\partial\mathbb{B}^{d}$. When $\gamma_{1},\gamma_{2}$ are two smooth conductivities that satisfy $\gamma_{1}|_{U_{s}}=\gamma_{2}|_{U_{s}}$ for some $s\in(0,1)$ it is known that $\Lambda_{\gamma_{1}}-\Lambda_{\gamma_{2}}$ is a pseudo-differential operator whose symbol (defined modulo smoothing operators) is zero: this implies that
$\lambda_{k}[\gamma_{1}]-\lambda_{k}[\gamma_{2}]=\mathcal{O}(k^{-\infty}),\qquad k\to\infty.$
A straightforward consequence of Theorems 1 and 2 is the following.
@@
This result extends in the Euclidean setting [DKN21, Theorem 1.4] to $d\geq 2$ and less regular conductivities. ˜2 has a quantitative counterpart, as we now show.
Define for $N>0$, $K>0$ and $1\leq p\leq\infty$ the set $\mathcal{G}_{K,B}^{p}$ consisting of those conductivities $\gamma\in W^{2,p}(\mathbb{B}^{d})$ that satisfy
(1.10) $K^{-1}\leq\gamma(x)\leq K,\quad x\in\mathbb{B},\qquad\left\lVert\gamma\right\rVert_{W^{2,p}(\mathbb{B}^{d})}\leq N.$
The Born approximation $\gamma^{\mathrm{B}}$ determines a conductivity $\gamma\in\mathcal{G}_{K,B}^{p}\subset{W^{2,1}}(\mathbb{B}^{d})$ in a Hölder continuous way.
Theorem 4 (Stability).
@@
˜4 shows that $\gamma$ can be determined by $\gamma^{\mathrm{B}}$ in a Hölder stable way, which is not the case for the problem of recovering a radial $\gamma$ from $\operatorname{\mathrm{Spec}}(\Lambda_{\gamma})$: this problem is ill-posed (no continuity unless the conductivities lie in a compact set [Ale88, AC08, FKR14] and the modulus of continuity is generally, at best, logarithmic [Man01]). Therefore, this shows that the ill-posedness of reconstructing a radial $\gamma$ from $\operatorname{\mathrm{Spec}}(\Lambda_{\gamma})$ is due to the linear step of solving the moment problem (1.4) or (1.6) to reconstruct $\gamma^{\mathrm{B}}$ (the moment problem is, indeed, a notoriously ill-posed problem). Thus, the introduction of the Born approximation $\gamma^{\mathrm{B}}$, which depends linearly on $\operatorname{\mathrm{Spec}}(\Lambda_{\gamma})$, allows transforming an ill-posed problem into a Hölder stable one.
One can compute the Born approximation of a conductivity explicitly in some special cases, see Section˜3.5; some of these explicit examples are shown in Fig.˜1. Also, it is possible to infer more precise information on the regularity of the Born approximation. In general, it is always a continuous function on $\mathbb{B}^{d}\setminus\{0\}$ that can be eventually singular at the origin. The nature of this singularity is related to the existence of eigenvalues and/or resonances for a certain 1-d Schrödinger operator defined in terms of $\gamma$, as will be shown in Section˜3.
Theorem 5 (Regularity).
@@
    $\gamma^{\mathrm{B}}(x)=-c_{0}\log|x|+\sum_{j=1}^{J}\frac{c_{j}}{|x|^{2\kappa_{j}}}+\gamma^{\mathrm{B}}_{\mathrm{reg}}(x),$
@@
From this point of view, the existence of the Born approximation proved in ˜1 provides a decomposition of the process of reconstructing $\gamma$ from $\Lambda_{\gamma}$ in two steps.
  • The first step (obtaining $\gamma^{\mathrm{B}}$ from $\Lambda_{\gamma}$) is linear and amounts to solving the moment problem (1.6), which is a notoriously ill-posed inverse problem.
  • The second step (obtaining $\gamma$ from $\gamma^{\mathrm{B}}$) is non-linear but enjoys good stability properties, as ˜4 shows.
The second property shows that the Born approximation induces a stable factorization of the Calderón problem in the radial case. Also, the local nature of $\gamma^{\mathrm{B}}$ provides a direct way to obtain local uniqueness for the radial Calderón problem (˜3) without using the standard approach based on Complex Geometrical Optics solutions of the equation (1.2). This approach based on the Born approximation has also been successfully applied to the Schrödinger operator version of the Calderón problem (the so-called Gelf’and-Calderón problem) at zero energy [DMMN24], and at fixed energy [MMSM25], and has been formulated in more general settings in [MM24]. We remark that, using the reconstruction algorithms in [DMMN24], one could transform ˜3 into an explicit reconstruction algorithm to determine $\gamma$ from $\Lambda_{\gamma}$.
@@ Acknowledgments
F.M. and C.M. have been supported by grant PID2021-124195NB-C31 from Agencia Estatal de Investigación (Spain). F. N. is supported by the GDR Dynqua.
2. Existence of the Born approximation
2.1. Linearized DtN operators in the radial case
Consider the non-linear map
(2.1) $\begin{array}[]{cccl}\Psi:&X&\longrightarrow&\mathcal{L}(H^{1/2}(\partial\mathbb{B}^{d}),H^{-1/2}(\partial\mathbb{B}^{d}))\\ &\gamma&\longmapsto&\Lambda_{\gamma}\end{array}$
where $X$ denotes the set of conductivities that are in $L^{\infty}(\mathbb{B}^{d})$. In this section, we obtain an explicit formula for the Fréchet differential in the radial case.
Proposition 2.1.
The Fréchet derivative $d\Psi_{\mathds{1}}$ of $\Psi$ in the space $L^{\infty}(\mathbb{B}^{d},\mathbb{R})$ at the conductivity $\gamma=\mathds{1}$ exists and it is a bounded operator
$d\Psi_{\mathds{1}}:L^{\infty}(\mathbb{B}^{d},\mathbb{R})\to\mathcal{L}(H^{1}(\mathbb{S}^{d-1}),L^{2}(\mathbb{S}^{d-1})).$
If, in addition, $h\in L^{\infty}_{\mathrm{rad}}(\mathbb{B}^{d})$, then $d\Psi_{\mathds{1}}(h)$ is the unbounded operator on $L^{2}(\mathbb{S}^{d-1})$, with domain $H^{1}(\mathbb{S}^{d-1})$, characterized by
(2.2) $d\Psi_{\mathds{1}}(h)|_{\mathfrak{H}_{k}}=k\left(2k+d-2\right)\sigma_{k-1}[h]\operatorname{\mathrm{Id}}_{\mathfrak{H}_{k}},\qquad k\in\mathbb{N},$
and $d\Psi_{\mathds{1}}(h)(1)=0$. In particular $\Lambda_{\mathds{1}}=d\Psi_{\mathds{1}}(\mathds{1})$, and $d\Psi_{\mathds{1}}$ admits a natural extension to $L^{1}_{\mathrm{rad}}(\mathbb{B}^{d})$ via identity (2.2).
Proof of Proposition˜2.1.
The Fréchet differential of $\Psi:\gamma\longmapsto\Lambda_{\gamma}$ in $L^{\infty}_{\mathrm{rad}}(\mathbb{B}^{d})$ at the constant conductivity $\gamma=1$ is given by the expression
$(f_{1},d\Psi_{\mathds{1}}(h)f_{2})_{L^{2}(\mathbb{S}^{d-1})}=\int_{\mathbb{B}^{d}}h(x)\overline{\nabla u_{1}(x)}\cdot\nabla u_{2}(x)\,dx,$
where $\Delta u_{j}=0$ in $\mathbb{B}^{d}$ and $u_{j}|_{\mathbb{S}^{d-1}}=f_{j}$, for $j=1,2$. Therefore, $d\Psi_{\mathds{1}}(h)$ is diagonalizable on the basis of spherical harmonics and its eigenvalues are:
(2.3) $(Y_{k},d\Psi_{\mathds{1}}(h)Y_{k})_{L^{2}(\mathbb{S}^{d-1})}=\int_{\mathbb{B}^{d}}h(x)\overline{\nabla Y_{k}^{s}(x)}\cdot\nabla Y_{k}^{s}(x)\,dx,$
where $Y_{k}^{s}(x):=|x|^{k}Y_{k}(x/|x|)$ stands for the solid harmonic corresponding to a normalized spherical harmonic $Y_{k}\in\mathfrak{H}_{k}$. When $k=0$, the solid harmonics are constant functions and (2.3) implies $d\Psi_{\mathds{1}}(h)(1)=0$.
@@ 2.1. Linearized DtN operators in the radial case
Direct computation gives that
$\nabla Y_{k}^{s}(x)=|x|^{k-1}\left(kY_{k}(\hat{x})\hat{x}+\nabla_{\mathbb{S}^{d-1}}Y_{k}(\hat{x})\right),\qquad\hat{x}=x/|x|.$
where $\nabla_{\mathbb{S}^{d-1}}$ denotes the gradient in $\mathbb{S}^{d-1}$. Writing $h=h_{0}(|\cdot|)$ and using that $\Delta_{\mathbb{S}^{d-1}}Y_{k}=k(k+d-2)Y_{k}$. $\left\lVert Y_{k}\right\rVert_{L^{2}(\mathbb{S}^{d-1})}=1$, one gets that (2.3) equals
$\left(k^{2}+\int_{\mathbb{S}^{d-1}}|\nabla_{\mathbb{S}^{d-1}}Y_{k}(x)|^{2}\,dS(x)\right)\int_{0}^{1}h_{0}(r)r^{2k-2+d-1}\,dr=k(2k-d-2)\sigma_{k-1}[h],$
as claimed. In the particular case $h=\mathds{1}$ this gives
$(Y_{k},d\Psi_{\mathds{1}}(\mathds{1})Y_{k})_{L^{2}(\mathbb{S}^{d-1})}=k$
which implies that $d\Psi_{\mathds{1}}(\mathds{1})=\Lambda_{\mathds{1}}$. ∎
2.2. The Born approximation for Schrödinger operator DtN maps
It is well-known that, provided $\gamma$ has enough regularity, the study of $\Lambda_{\gamma}$ can be reduced to that of a DtN map associated to a certain Schrödinger operator. If $\gamma\in W^{2,p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R}_{+})$ with $p>d/2$ is a conductivity, and $u$ solves (1.2) then $v=\sqrt{\gamma}u$ satisfies
(2.4) $\begin{cases}-\Delta v+Vv&=0,\qquad\text{on }\mathbb{B}^{d},\\ \hfill v|_{\mathbb{S}^{d-1}}&=\sqrt{\gamma}f,\end{cases}$
where
(2.5) $V:=\frac{\Delta\sqrt{\gamma}}{\sqrt{\gamma}}\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R}).$
One can check that the following identity holds
(2.6) $\Lambda^{\mathrm{p}}_{V}f:=\partial_{\nu}v|_{\mathbb{S}^{d-1}}=\gamma^{-1/2}\Lambda_{\gamma}(\gamma^{-1/2}f)+\frac{1}{2}\gamma^{-1}(\partial_{\nu}\gamma)f,$
when $v$ solves (2.4) with $V$ given by (2.5). From this identity, it is clear that the Calderón problem for conductivities is equivalent to the one with Schrödinger operators provided $\gamma$ and $\partial_{\nu}\gamma$ are known at the boundary $\mathbb{S}^{d-1}$.
Again, since (2.4) is invariant by rotations,
(2.7) $\Lambda^{\mathrm{p}}_{V}|_{\mathfrak{H}_{k}}=\lambda^{\mathrm{p}}_{k}[V]\operatorname{\mathrm{Id}}_{\mathfrak{H}_{k}},\qquad\forall k\in\mathbb{N}_{0}.$
Moreover, when $V=0$ a direct computation gives that
$\Lambda^{\mathrm{p}}_{0}(Y_{k})=kY_{k}\qquad\forall Y_{k}\in\mathfrak{H}_{k},$
and hence, $\lambda^{\mathrm{p}}_{k}[0]=\lambda_{k}[\mathds{1}]=k$ for every $k\in\mathbb{N}_{0}$, since one has $\Lambda^{\mathrm{p}}_{0}=\Lambda_{\mathds{1}}$ from (2.6).

### v2

@@ Abstract.
The problem of characterizing sequences of real numbers that arise as spectra of Dirichlet-to-Neumann (DtN) maps for elliptic operators has attracted considerable attention over the past fifty years. In this article, we address this question in the simple setting of DtN maps associated with a rotation-invariant elliptic operator $\operatorname{div}(\gamma\nabla\cdot)$ in the ball in Euclidean space. We show that the spectrum of such a DtN operator can be expressed as a universal term, determined solely by the boundary values of the conductivity $\gamma$, plus a sequence of Hausdorff moments of an integrable function, which we call the Born approximation of $\gamma$. We also show that this object is locally determined from the boundary by the corresponding values of the conductivity, a property that implies a local uniqueness result for the Calderón Problem in this setting. We also give a stability result: the functional mapping the Born approximation to its conductivity is Hölder stable in suitable Sobolev spaces. Finally, in order to refine the characterization of the Born approximation, we analyze its regularity properties and their dependence on the conductivity.
@@ 1. Introduction
where $\nu$ is the outward normal unit vector field on $\partial\Omega$, $\partial_{\nu}=\nu\cdot\nabla$ and $u$ is the solution to the conductivity equation
@@
Since constants are always solutions to (1.2), one always has $\lambda_{0}[\gamma]=0$. Also, if $\gamma$ is identically $1$ in $\mathbb{B}^{d}$, one has $\lambda_{k}[\gamma]=k$, for all $k\in\mathbb{N}_{0}$. In general, we will show that the rest of the eigenvalues of a radial DtN map have a very rigid structure: they can be expressed in terms of moments of radial functions. For $g\in L^{1}(\mathbb{B}^{d})$ define
(1.3) $\sigma_{k}[g]:=\frac{1}{|\mathbb{S}^{d-1}|}\int_{\mathbb{B}^{d}}g(x)|x|^{2k}\,dx,\qquad k\in\mathbb{N}_{0},$
@@
(1.4) $\lambda_{k}[\gamma]=\gamma|_{\mathbb{S}^{d-1}}k-\frac{1}{2}\partial_{\nu}\gamma|_{\mathbb{S}^{d-1}}+\frac{1}{2}\sigma_{k}[\Delta\gamma^{\mathrm{B}}],\qquad\forall k\in\mathbb{N},$
@@
(1.5) $\gamma^{\mathrm{B}}|_{\mathbb{S}^{d-1}}=\gamma|_{\mathbb{S}^{d-1}},\qquad\qquad\partial_{\nu}\gamma^{\mathrm{B}}|_{\mathbb{S}^{d-1}}=\partial_{\nu}\gamma|_{\mathbb{S}^{d-1}}.$
@@
    (1.6) $\sigma_{k}[\gamma^{\mathrm{B}}]=\frac{\lambda_{k+1}[\gamma]}{2(k+1)(k+1+\nu_{d})},\qquad k\in\mathbb{N}_{0},\qquad\nu_{d}:=\frac{d-2}{2}.$
@@
$\sigma_{k}[\Delta\gamma^{\mathrm{B}}]=\int_{0}^{1}h_{\gamma}(t)t^{k}\,dt,\qquad k\in\mathbb{N}_{0}.$
@@
The function $\gamma^{\mathrm{B}}$ appearing in ˜1 is called the Born approximation of $\gamma$. It is a purely spectral object, in the sense that it is completely determined by $\operatorname{\mathrm{Spec}}(\Lambda_{\gamma})$ as (1.7) shows. However, it is local in the sense that, in any annular neighborhood of the boundary $\partial\mathbb{B}^{d}$, it is uniquely determined by the values of the conductivity in that same neighborhood.
@@
Structural results for the spectra of DtN maps corresponding to general, smooth, non-radial conductivities can be obtained by a detailed analysis of (pseudodifferential or Berezin) symbols of DtN maps; see, for instance, [LSA21, Roz78] for the two-dimensional case or [PEUVB24] for the ball in Euclidean space. However, if $\gamma_{1},\gamma_{2}$ are two smooth conductivities that satisfy $\gamma_{1}|_{U_{s}}=\gamma_{2}|_{U_{s}}$ for some $s\in(0,1)$ it is known that $\Lambda_{\gamma_{1}}-\Lambda_{\gamma_{2}}$ is a pseudo-differential operator whose symbol –defined modulo smoothing operators– is zero (in the radial context, this implies that $\lambda_{k}[\gamma_{1}]-\lambda_{k}[\gamma_{2}]=\mathcal{O}(k^{-\infty})$ as $k\to\infty$). Therefore, these approaches are unable to determine how the interior behavior of $\gamma$ affects the spectrum. In contrast, we observe that ˜1 provides a non-trivial description of the spectrum, even in the case where the conductivity $\gamma$ is constant in a neighborhood of the boundary $\partial\mathbb{B}^{d}$. In fact, a straightforward consequence of Theorems 1 and 2 is the following.
@@
This result extends [DKN21, Theorem 1.4] to $d\geq 2$ and less regular conductivities in the Euclidean setting. Due to its local nature, this result is stronger than particularizing to the radial case the standard uniqueness results for the Calderón problem. We also notice that the proof of the corollary and the previous theorems does not involve the standard approach based on Complex Geometrical Optics solutions of the equation (1.2). An analogous result for the case of DtN maps associated to radial Schrödinger operators is given in Theorem˜3.12 in Section˜3.6.
˜2 has a quantitative counterpart, as we now show. Define for $N>0$, $K>0$ and $1\leq p\leq\infty$ the set $\mathcal{G}_{K,N}^{p}$ consisting of those conductivities $\gamma\in W^{2,p}_{\mathrm{rad}}(\mathbb{B}^{d})$ that satisfy
(1.10) $K^{-1}\leq\gamma(x)\leq K,\quad x\in\mathbb{B}^{d},\qquad\left\lVert\gamma\right\rVert_{W^{2,p}(\mathbb{B}^{d})}\leq N.$
The Born approximation $\gamma^{\mathrm{B}}$ determines a conductivity $\gamma\in\mathcal{G}_{K,N}^{p}\subset{W^{2,1}}(\mathbb{B}^{d})$ in a Hölder continuous way.
Theorem 4.
@@
˜4 shows that $\gamma$ can be determined by $\gamma^{\mathrm{B}}$ in a Hölder stable way, which is not the case for the problem of recovering a radial $\gamma$ from $\operatorname{\mathrm{Spec}}(\Lambda_{\gamma})$: this problem is ill-posed (no continuity unless the conductivities lie in a compact set [Ale88, AC08, FKR14] and the modulus of continuity is generally, at best, logarithmic [Man01]). Therefore, this shows that the ill-posedness of reconstructing a radial $\gamma$ from $\operatorname{\mathrm{Spec}}(\Lambda_{\gamma})$ is due to the linear step of solving the moment problem (1.4) or (1.6) to reconstruct $\gamma^{\mathrm{B}}$ (the moment problem is, indeed, a notoriously ill-posed problem).
This shows that the existence of the Born approximation proved in ˜1 provides a decomposition of the process of reconstructing $\gamma$ from $\Lambda_{\gamma}$ in two steps:
  • The first step (obtaining $\gamma^{\mathrm{B}}$ from $\Lambda_{\gamma}$) is linear and amounts to solving the moment problem (1.6).
  • The second step (obtaining $\gamma$ from $\gamma^{\mathrm{B}}$) is non-linear but enjoys good stability and locality properties, as ˜4 shows.
In other words, the existence of the Born approximation and this decomposition provide a stable factorization of the Calderón problem in the radial case.
It is possible to infer more precise information on the regularity of the Born approximation. In general, it is always a continuous function on $\mathbb{B}^{d}\setminus\{0\}$ that can be eventually singular at the origin. The nature of this singularity is related to the existence of eigenvalues and/or resonances for a certain 1-d Schrödinger operator defined in terms of $\gamma$, as will be shown in Section˜3.
Theorem 5.
@@
    $\gamma^{\mathrm{B}}(x)=\gamma^{\mathrm{B}}_{\mathrm{reg}}(x)-c_{0}\log|x|+\sum_{j=1}^{J}\frac{c_{j}}{|x|^{2\kappa_{j}}},$
@@
One can compute the Born approximation of a conductivity explicitly in some special cases, see Section˜3.5; some of these explicit examples are shown in Fig.˜1 and Fig.˜2. These are chosen to illustrate that $\gamma^{\mathrm{B}}$ can have a logarithmic singularity when $d=3$ (Fig.˜1), and to illustrate the continuity of the Born approximation (Fig.˜2) when $d=2$ (and $d=3$ in some special cases).
@@
By analogy with scattering problems, the name of Born approximation is motivated by the fact that $\gamma^{\mathrm{B}}$ is a linearization of the Calderón inverse problem, as (1.6) shows ($\gamma^{\mathrm{B}}$ depends linearly on the spectrum). Similarly as was explained in detail in [DMMN24] for the case of DtN maps associated to Schrödinger operators, or more systematically in [CMMSM26], one can show, as a consequence of ˜1, that $\gamma^{\mathrm{B}}$ satisfies the identity
(1.11) $\Lambda_{\gamma}=\mathrm{d}\Psi_{1}(\gamma^{\mathrm{B}}),$
where $\Psi:\gamma\longmapsto\Lambda_{\gamma}$ and $\mathrm{d}\Psi_{1}$ denotes the Fréchet differential of $\Psi$ at the constant conductivity $\gamma=1$. As far as we know this is the first time in the literature that (1.11) has been shown to have an exact solution. This is, in fact, a long standing open question in the field of numerical approaches to the Calderón problem, see [CMMSM26] and [HS10] for more details.
Figure 1. Plot of the radial profile of $\gamma_{d,\mu,\nu}$ (blue) and $\gamma^{\mathrm{B}}_{d,\mu,\nu}$ (orange) defined in Section˜3.5 for $d=3,\;\mu=1,\;\nu=0$. One can observe the presence of a logarithmic singularity in the Born approximation induced by a resonance at zero.
The approach based on the Born approximation has also been successfully applied in the radial case to the Schrödinger operator version of the Calderón problem (the so-called Gelf’and-Calderón problem) at zero energy [DMMN24], and at fixed energy [MMSM25], and has been formulated in more general settings in [MM24]. In particular, analogous results to Theorems 1-4 can be found in [DMMN24, MMSM25] for the respective problems. Linearization at other conductivities besides the constant conductivities and existence of the corresponding Born approximation has been addressed in [CMMSM26], as well as some partial results for the non-radial case in two dimensions. We remark that, using the reconstruction algorithms in [DMMN24], one could transform ˜3 into an explicit reconstruction algorithm to determine $\gamma$ from $\Lambda_{\gamma}$.
@@ Acknowledgments
F.M. and C.M. have been supported by grants PID2021-124195NB-C31 and PID2024-158664NB-C21 from Agencia Estatal de Investigación (Spain). F. N. is supported by the GDR Dynqua.
@@
2. Existence of the Born approximation
2.1. The Born approximation for DtN maps of Schrödinger operators
It is well-known that, provided $\gamma$ has enough regularity, the study of $\Lambda_{\gamma}$ can be reduced to that of a DtN map associated with a certain Schrödinger operator. If $V\in L^{p}(\mathbb{B}^{d},\mathbb{R})$ with $p>d/2$ is such that zero is not a Dirichlet eigenvalue of $\Delta+V$ on $L^{2}(\mathbb{B}^{d})$, then for every $f\in H^{1/2}(\mathbb{S}^{d-1})$, the problem
(2.1) $\begin{cases}-\Delta v+Vv&=0,\qquad\text{on }\mathbb{B}^{d},\\ \hfill v|_{\mathbb{S}^{d-1}}&=f,\end{cases}$
admits a unique solution $v\in H^{1}(\mathbb{B}^{d})$. The DtN map is defined by the identity $\Lambda^{\mathrm{S}}_{V}f=\partial_{\nu}v|_{\mathbb{S}^{d-1}}$, where $v$ is the solution of (2.1). If the regularity of $V$ or $f$ is not enough to define $\Lambda^{\mathrm{S}}_{V}f$, one uses the standard weak definition:
$\left\langle g,\Lambda^{\mathrm{S}}_{V}f\right\rangle_{H^{1/2}(\mathbb{S}^{d-1})\times H^{-1/2}(\mathbb{S}^{d-1})}:=\int_{\mathbb{B}^{d}}\left(\nabla u\cdot\nabla v+Vuv\right)\,dx,$
where $v$ solves (2.1) and $u$ is any $H^{1}(\mathbb{B}^{d})$ function with trace $g$.
If $\gamma\in W^{2,p}(\mathbb{B}^{d},\mathbb{R}_{+})$ with $p>d/2$ is a conductivity and $u$ solves (1.2), then $v=\sqrt{\gamma}u$ satisfies (2.1) with boundary datum $\sqrt{\gamma}f$, where
(2.2) $V:=\frac{\Delta\sqrt{\gamma}}{\sqrt{\gamma}}\in L^{p}(\mathbb{B}^{d},\mathbb{R}).$
Moreover, if $V$ arises in this way from a conductivity, then zero is not a Dirichlet eigenvalue of $\Delta+V$ on $L^{2}(\mathbb{B}^{d})$, and one can check that the following identity holds
(2.3) $\Lambda^{\mathrm{S}}_{V}f=\gamma^{-1/2}\Lambda_{\gamma}(\gamma^{-1/2}f)+\frac{1}{2}\gamma^{-1}(\partial_{\nu}\gamma)f.$
From this identity, it is clear that the Calderón problem for conductivities is equivalent to the one with Schrödinger operators, provided $\gamma$ and $\partial_{\nu}\gamma$ are known at the boundary $\mathbb{S}^{d-1}$.
As was the case with conductivities, if $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$, then $\Lambda^{\mathrm{S}}_{V}$ is a self-adjoint unbounded operator on $L^{2}(\mathbb{S}^{d-1})$ that is invariant by rotations (since the equation (2.1) is). Therefore, it is diagonalizable in the basis of spherical harmonics. We denote:
$\Lambda^{\mathrm{S}}_{V}|_{\mathfrak{H}_{k,d}}=\lambda^{\mathrm{S}}_{k}[V]\operatorname{\mathrm{Id}}_{\mathfrak{H}_{k,d}},\qquad\forall k\in\mathbb{N}_{0}.$
Let $\mathds{1}\in L^{\infty}(\mathbb{B}^{d})$ be the conductivity that is identically 1 in $\mathbb{B}^{d}$. When $\gamma=\mathds{1}$, the associated potential is $V=0$ and hence $\Lambda^{\mathrm{S}}_{0}=\Lambda_{\mathds{1}}$ and $\lambda^{\mathrm{S}}_{k}[0]=\lambda_{k}[\mathds{1}]=k$ for every $k\in\mathbb{N}_{0}$.